Consider the following alternatives:

Year

Alternative A

Alternative B

Alternative C

0

-$400.00

-$500.00

$0.00

1

$163.00

$193.00

$0.00

2

$163.00

$193.00

$0.00

3

$163.00

$193.00

$0.00

4

$163.00

$193.00

$0.00

Given the following MARR rates, which alternative should be selected:

a) 6%

b) 9%

c) 10%

d) 15%
Answer #1

To see which alternative is preferable, we will calculate their present values

Present value = Future value/(1+Interest rate)^time

So for year 0 present value will be the same

For year 1 alternative A at 6% interest rate

= 163/(1+6%)^1

= $153.77

For the second year

= 163/(1+6%)^2

= $145.07

And same for Alternative B at different interest rates

At 9% interest rate, Alternative B

= 193/(1+9%)^1

= $177.06

Alternative C is do nothing, which will be preferred if the present value of both of them is negative.

The following tales will show the present value of both alternative

Year Alternative A PV Alternative B PV
0 -$400 -$400.00 -$500 -$500.00
1 $163 $153.77 $193 $182.08
2 $163 $145.07 $193 $171.77
3 $163 $136.86 $193 $162.05
4 $163 $129.11 $193 $152.87
Total $164.81 $168.77

Aletrnative B is preferrable

b) At interest rate of 9%

Year Alternative A PV Alternative B PV
0 -$400 -$400.00 -$500 -$500.00
1 $163 $149.54 $193 $177.06
2 $163 $137.19 $193 $162.44
3 $163 $125.87 $193 $149.03
4 $163 $115.47 $193 $136.73
Total $128.07 $125.27

Alternative A will be preferred.

c) At interest rate of 10%

Year Alternative A PV Alternative B PV
0 -$400 -$400.00 -$500 -$500.00
1 $163 $148.18 $193 $175.45
2 $163 $134.71 $193 $159.50
3 $163 $122.46 $193 $145.00
4 $163 $111.33 $193 $131.82
Total $116.69 $111.78

Alternative A will be preferred.

d) At interest rate of 15%

Year Alternative A PV Alternative B PV
0 -$400 -$400.00 -$500 -$500.00
1 $163 $141.74 $193 $167.83
2 $163 $123.25 $193 $145.94
3 $163 $107.18 $193 $126.90
4 $163 $93.20 $193 $110.35
Total $65.36 $51.01

Alternative A will be prferred.